# Analytic structure of stress-energy response functions and new Kubo formulae

### v1

@@ II.1 Ward Identity Analysis
0$\displaystyle$ 0 $\displaystyle=k_{\alpha}\left(\tilde{G}^{\alpha\beta,\mu\nu}(\omega,{\bf k}){-}\eta^{\beta\mu}\langle\hat{T}^{\alpha\nu}\rangle{-}\eta^{\beta\nu}\langle\hat{T}^{\alpha\mu}\rangle{+}\eta^{\alpha\beta}\langle\hat{T}^{\mu\nu}\rangle\right)$
@@
The flat-space metric we use is $\eta_{\mu\nu}=\hbox{diag}(-1,1,1,1)$. Here, $\langle{\hat{\cal{O}}}\rangle=(1/Z)\hbox{Tr}\,(e^{-\beta\hat{H}}\hat{\cal{O}})$ is the thermal average of the operator $\hat{\cal{O}}$. Using the CPT properties of the stress-energy tensor operators, one can show that $G_{R}^{\alpha\beta,\mu\nu}(\omega,{\bf k})=G_{R}^{\mu\nu,\alpha\beta}(\omega,{\bf k})$.
@@
Using the above Ward identities, all correlators involving the energy-momentum density operator $\hat{T}^{0\mu}$ can be expressed in terms of $\tilde{G}^{{ij},{lm}}(\omega,{\bf k})$. In a spatially isotropic system, this correlator can be decomposed in terms of the five orthogonal rank-2 tensors involving $\delta_{ij}=\eta^{ij}$ and $k_{i}k_{j}$ [34] as follows 222 This expression first appeared in an unpublished note by L.G. Yaffe.
@@ V Third order conformal hydrodynamics and the relaxation times
where $\varphi_{-1|0}=8T^{4}/5\pi^{2}$, and $\xi^{lij}$ and $\zeta^{ijkl}$ are the new spin-3 and spin-4 degrees of freedom, respectively. These are still conformal because $\partial_{l}\xi^{lij}\sim\partial_{l}\partial^{\langle l}\pi^{ij\rangle}$ can be decomposed as a linear combination of $\ij\displaystyle\nabla^{2}\pi^{\ij}$ and $\partial^{\langle i}\partial_{l}\pi^{lj\rangle}$. Both are Weyl invariant [69]. One can also get the same set of linearized equations from [62] (up to $\xi^{lij}$), and a slightly different set from de Brito and Denicol [59] (up to $\zeta^{ijkl}$) that contains slightly different coefficients. The differences come from three different methods employed in these studies to get $\xi^{lij}$ and $\zeta^{ijkl}$. As this difference does not change the qualitative conclusion of this section, we will only consider the above equations.
@@ VI New Kubo Formulae
Only $G_{2}$ and $G_{T}$ contribute to this Kubo formula since all others vanish in the $\omega\to 0$ limit with $k\neq 0$. Note that deriving Eq.(72) is possible only because we have the analytic structures of all five response functions. Note that even though both Eq.(71) and Eq.(72) can be used to calculate $\eta$, the two orders of the limits do not commute.
@@
We can derive a few more Kubo formulae for $\eta$ and $\zeta$ all with the $\omega\to 0$ limit taken first. Noting $G_{R}^{xx,xx}(\omega,k_{z})=P+G_{2}+G_{T}$, we get
$\displaystyle\lim_{k_{z}\to 0}\lim_{\omega\to 0}\partial_{\omega}G^{xx,xx}_{R}(\omega,k_{z})$ $\displaystyle=-4i\eta$ (77)
@@
$\displaystyle\lim_{\omega\to 0}\lim_{k_{z}\to 0}\partial_{\omega}G^{xx,xx}_{R}(\omega,k_{z})$ $\displaystyle=-i\left(\zeta+{4\eta\over 3}\right)$ (78)
@@ VII On diagrammatic calculations of viscosities
$\displaystyle{\rm Im}G_{R}^{xy,xy}(\omega,k_{z})$ $\displaystyle={1\over 2}(1-e^{-\beta\omega})\sigma^{xy,xy}(\omega,k_{z})$ (88)
where $\sigma^{xy,xy}(\omega,k_{z})$ is the Wightman function, we only need to examine
$\displaystyle\sigma^{xy,xy}_{\hbox{\scriptsize 1-loop}}(\omega,k_{z})$ $\displaystyle=2\int{d^{4}p\over(2\pi)^{4}}\,p_{x}^{2}p_{y}^{2}n(p^{0})(1+n(p^{0}))$
@@
$\displaystyle=\int{d^{4}p\over(2\pi)^{4}}\,n(p^{0})(1+n(p^{0}))$
@@ Appendix B Correlation functions in terms of the response functions
$\displaystyle\omega^{3}G_{R}^{00,0i}(\omega,k)$ $\displaystyle=k_{i}k^{2}\left(G_{L}(\omega,k^{2})-{\omega^{2}\over k^{2}}h_{0}\right)$ (99)
$\displaystyle\omega^{4}G_{R}^{00,00}(\omega,k)$ $\displaystyle=k^{4}\left(G_{L}(\omega,k^{2})-{\omega^{2}\over k^{2}}h_{0}\right)$ (100)

### v2

@@ II.1 Ward Identity Analysis
$\displaystyle 0$ $\displaystyle=k_{\alpha}\left(\tilde{G}^{\alpha\beta,\mu\nu}(\omega,{\bf k}){-}\eta^{\beta\mu}\langle\hat{T}^{\alpha\nu}\rangle{-}\eta^{\beta\nu}\langle\hat{T}^{\alpha\mu}\rangle{+}\eta^{\alpha\beta}\langle\hat{T}^{\mu\nu}\rangle\right)$
@@
[…]
@@
Using the above Ward identities, all correlators involving the energy-momentum density operator $\hat{T}^{0\mu}$ can be expressed in terms of $\tilde{G}^{{ij},{lm}}(\omega,{\bf k})$. In a spatially isotropic system, this correlator can be decomposed in terms of the five orthogonal rank-2 tensors involving $\delta_{ij}=\eta^{ij}$ and $k_{i}k_{j}$ [34] as follows  [^2] This expression first appeared in an unpublished note by L.G. Yaffe.
@@ V Third order conformal hydrodynamics and the relaxation times
where $\varphi_{-1|0}=8T^{4}/5\pi^{2}$, and $\xi^{lij}$ and $\zeta^{ijkl}$ are the new spin-3 and spin-4 degrees of freedom, respectively. These are still conformal because $\partial_{l}\xi^{lij}\sim\partial_{l}\partial^{\langle l}\pi^{ij\rangle}$ can be decomposed as a linear combination of $\nabla^{2}\pi^{\penalty 10000\thinspace i\kern-0.16391pt\penalty 10000\thinspace j}$ and $\partial^{\langle i}\partial_{l}\pi^{lj\rangle}$. Both are Weyl invariant [69]. One can also get the same set of linearized equations from [62] (up to $\xi^{lij}$), and a slightly different set from [59] (up to $\zeta^{ijkl}$) that contains slightly different coefficients. The differences come from three different methods employed in these studies to get $\xi^{lij}$ and $\zeta^{ijkl}$. As this difference does not change the qualitative conclusion of this section, we will only consider the above equations.
@@ VI New Kubo Formulae
Only $G_{2}$ and $G_{T}$ contribute to this Kubo formula since all others vanish in the $\omega\to 0$ limit with $k\neq 0$. Note that deriving Eq.(72) is possible only because we have the analytic structures of all five response functions. Note also that even though both Eq.(71) and Eq.(72) can be used to calculate $\eta$, the two orders of the limits do not commute.
@@
We can derive a few more Kubo formulae for $\eta$ and $\zeta$ all with the $\omega\to 0$ limit taken first. Noting $\tilde{G}^{xx,xx}(\omega,k_{z})=P+G_{2}+G_{T}$, we get
$\displaystyle\lim_{k_{z}\to 0}\lim_{\omega\to 0}\partial_{\omega}\tilde{G}^{xx,xx}(\omega,k_{z})$ $\displaystyle=-4i\eta$ (77)
@@
$\displaystyle\lim_{\omega\to 0}\lim_{k_{z}\to 0}\partial_{\omega}\tilde{G}^{xx,xx}(\omega,k_{z})$ $\displaystyle=-i\left(\zeta+{4\eta\over 3}\right)$ (78)
@@ VII On diagrammatic calculations of viscosities
$\displaystyle{\rm Im}G_{R}^{xy,xy}(\omega,k_{z})$ $\displaystyle={1\over 2}(1-e^{-\beta\omega})\sigma_{\color[rgb]{1,0,0}>}^{xy,xy}(\omega,k_{z})$ (88)
where $\sigma_{\color[rgb]{1,0,0}>}^{xy,xy}(\omega,k_{z})$ is the greater Wightman function, we only need to examine
$\displaystyle\sigma^{xy,xy}_{{\color[rgb]{1,0,0}>,}\hbox{\scriptsize 1-loop}}(\omega,k_{z})$ $\displaystyle=2\int{d^{4}p\over(2\pi)^{4}}\,p_{x}^{2}p_{y}^{2}n(p^{0}{\color[rgb]{1,0,0}+\omega})(1+n(p^{0}))$
@@
$\displaystyle=\int{d^{4}p\over(2\pi)^{4}}\,n(p^{0}{{\color[rgb]{1,0,0}+\omega}})(1+n(p^{0}))$
@@ Appendix B Correlation functions in terms of the response functions
$\displaystyle\omega^{3}G_{R}^{00,0i}(\omega,k)$ $\displaystyle=k_{i}k^{2}\left(G_{L}(\omega,k^{2})+{\omega^{2}\over k^{2}}h_{0}\right)$ (99)
$\displaystyle\omega^{4}G_{R}^{00,00}(\omega,k)$ $\displaystyle=k^{4}\left(G_{L}(\omega,k^{2})+{\omega^{2}\over k^{2}}h_{0}\right)$ (100)